OpenStudy (anonymous):

3.Write the equation of the line passing through the points (0, -4) and (-2, 2).

OpenStudy (anonymous):

First find the Slope of the two given points. \[m = \frac{ y2 - y1 }{ x2 - x1 }\]\[m = \frac{ 2 - (-4) }{ -2 - 0} = \frac{ 2 + 4 }{ -2 }=\frac{ 6 }{ -2 }= -3\]

OpenStudy (anonymous):

Do you know the Point-Slope form?

OpenStudy (anonymous):

yes y-y1m(x-x1)?

OpenStudy (anonymous):

Right. So we substitute the values of a point, lets just use (0, -4), and slope into the equation. y - y1 = m(x - x1)

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

idk how to set it up

OpenStudy (anonymous):

y1 and x1 stand for the values of the coordinates, (x, y). (0, -4) 0 is the x-value, or x1 -4 is the y-value, or y1 Remember that our slope, m, is -3. So we substitute these into the equation: \[y - y1 = m(x - x1)\]\[y - (-4) = -3(x - 0)\]^Two negatives = a positive. \[y + 4 = -3(x - 0)\]Do you understand so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great! Now, I'm not sure if you need to convert it into Slope-Intercept form. Do you have any answer choices?

OpenStudy (anonymous):

@Jessiesummer do you know what format you need the answer to be in? Slope-Intercept: y = mx + b or Point-Slope: y - y1 = m(x - x1)